RawlsK29:

Consider the following pair of equations: y = x + 4 y = −2x − 2 Explain how you will solve the pair of equations by substitution. Show all the steps and write the solution in (x, y) form.

RawlsK29:

@EndersWorld

dude:

You have to set them equal to each other since they both equal y \(x + 4 =−2x − 2\)

RawlsK29:

Then subtract the x?

dude:

Yes you can do that

BenLindquist:

y = x + 4; y = −2x − 2 Solve y = x + 4 for y: Substitute x + 4 for y in y = −2x − 2: y = −2x − 2 x + 4 = −2x − 2 Add 2x to both sides x + 4 + 2x = −2x − 2 + 2x 3x + 4 = −2 Add -4 to both sides 3x + 4 + −4 = −2 + −4 3x = −6 Divide both sides by 3 3x/3 = −6/3 x = −2 Substitute −2 for x in y = x + 4: y = x + 4 y = −2 + 4 Simplify both sides of the equation y = 2 Answer: x = −2 and y = 2

RawlsK29:

Oh okay, thanks! Sorry @dude I was afk